OpenStudy (anonymous):

rewrite the expression (1/1-sinx)-(sinx/1+sinx)

OpenStudy (anonymous):

\[\frac{ 1 }{ 1-\sin x }-\frac{ \sin x }{ 1+\sin x } \] =\[\frac{ 1+\sin x-\sin x(1-\sin x) }{ 1-\sin ^2x }\] =\[\frac{ 1+\sin x-\sin x+\sin ^2x }{ \cos ^2x }=\frac{ 1+\sin ^2x }{ \cos ^2x }=\frac{ -\cos ^2x }{ \cos ^2x }\] =-1

OpenStudy (anonymous):

thank you